OpenStudy (anonymous):

14 1/3 + 5 4/9?? simplify.

OpenStudy (anonymous):

14 1/3 has whole part 14 , fraction 1/3 5 4/9 has whole part 5 and fraction 1/9 14 1/3 + 5 4/9 = 14 +5 + 1/3 +1/9 = 19 + [ 1/3 +1/9] Least common denominator is 9 1/3 +1/9 = 3/9 +1/9 = 4/9 Answer is 19 4/9